Fermi National Laboratory


Accelerator Update

Monday May 17
The day shift began with Operations monitoring store 3504, with Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

At 8:43 AM, the Collider Detector at Fermilab (CDF) called asking that the colliding beam problems get resolved so they could turn their detector on. Operations called Tevatron (TeV) experts who made adjustments and had CDF back on line forty-five minutes later.

TeV experts began end-of-store studies at 10:20 AM.

Pbar experts began a stacking study at 10:40 AM, followed by a reverse proton study around noon.

Several Klystron vacuum valves closed at 11:34 AM. Operators investigated but couldn't find a problem. After tripping twice more, operators discovered welders working in the area and whenever they struck an arc it caused the trip. The work was halted. All the valves opened except for one. Operations prepared the Linac tunnel for an access at 12:20 PM. Linac experts power cycled the valve controller, which solved the problem around 1:43 PM.

The end-of-store studies ended at 11:39 AM and Operations terminated store 3504. Operators prepared the TeV tunnels and the CDF and D-zero (D0) collision halls for access at 11:52 AM.

A Main Injector (MI) beam valve closed at 2 PM and wouldn't reopen. Operations called in experts to investigate. At 2:30 PM, Operations noticed that the beam valve had opened, but no one had reported doing anything.

Operations resumed sending beam to MiniBooNE at 2:04 PM.

Operations received a report at 2:27 PM that the lighting was off in TeV sector E0. Operators discovered other systems off due to a local power outage. The Duty Electrician found and reset a tripped breaker at E0.

D0 came out of access at 4:03 PM.

Operations established beam to MTest at 4:09 PM and to MCenter at 4:36 PM.

MTest finished using beam at 6:31 PM.

Pbar studies ended at 7:11 PM and stacking resumed.

Operations put the TeV into shot setup at 10:24 PM. Beam to Switchyard (SY) was terminated, as always, during the setup.

Recycler experts completed their studies at 10:36 PM.

Tuesday May 18
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

At 12:47 AM, Operations established store 3506 with an initial luminosity of 60.31E30. Pbar resumed stacking.

Operations reported at 2:25 AM that there was a colliding beam problem keeping off both CDF and D0. Operators made collimator adjustments, but the problem remained. They kept working on the problem

Operations established beam to SY and the MCenter experiment at 2:32 AM.

CDF and D0 were both able to turn the detector high voltages back on and resume taking data at 3:46 AM.

A Recycler expert began a study at 5 AM.

At 5:48 AM, Operations reported that the TeV sector C3 cold compressor tripped off and caused a quench. Cryo system technicians investigated.

MCenter finished using beam at 7:20 AM.

Cryo system techs reported at 8:23 AM that the C3 cold compressor checked out okay.

Operation put the TeV through a dry squeeze at 11:18 AM. At the end of this squeeze, TeV experts will then conduct study-related wet squeezes.

A Linac controls expert began rebooting some nodes at 3:22 PM. This halted all downstream beam until 4:18 PM. Operations resumed sending beam to MiniBooNE.

A NuMI expert began a study at 5 PM and ended the study at 6:00 PM.

A NTF expert began an hour and a half study at 6:33 PM.

At 7:53 PM, Operations established store 3510 with an initial luminosity of 62.76E30. Pbar resumed stacking.

Wednesday May 19
The midnight shift began with Operations monitoring store 3510, with Pbar stacking, and with MiniBooNE taking beam.

At 2:08 AM, CDF called about a problem with the colliding beam. Operators tuned the beam a little and fixed the problem. At 4:08 AM, the problem returned. Operators tried the same fix, but it didn't work. A TeV expert suggested that operators turn off the TeV dampers, which they did.

At 5:17 AM, the TeV colliding beam problem returned. Operators investigated.

A Recycler expert began a study at 6:34 AM.

A Booster vacuum burst caused a vacuum pump to trip off at 6:42 AM. It reset okay.

Plans for Today
The plans for to day include TeV and Pbar studies along with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/19/2004   email Fermilab